Circular No. 6945

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cd IN ANONYMOUS GALAXY A. V. Filippenko, D. C. Leonard, and M. Modjaz, University of Cali-

fornia at Berkeley; and R. G. Eastman, Lawrence Livermore National Laboratory, report that CCD spectra (range 320–1000 nm) obtained on June 18 UT with the Shane 3-m reflector at Lick Observatory show that SN 1998cd (cf. IAUC 6909) is of type Ia, roughly 6 weeks past maximum brightness.

1998 KY_{26}

M. Hicks and D. Rabinowitz, Jet Propulsion Laboratory, report: “We obtained BVRI photometry of this object with the Table Mountain 0.6m telescope on the night of June 3. After subtraction of solar colors, we determine color ratios of B − R = +0.050 ± 0.061, V − R = +0.024 ± 0.040, I − R = +0.045 ± 0.040, R = 17.08 ± 0.020, consistent with C- and M-type asteroids. R-band photometry from June 8.329–8.487 UT yields a double-peaked lightcurve with a mean absolute Bessel R magnitude of 24.95 ± 0.01 and a synodic period of 10.686 ± 0.024 min, consistent with the small size and extremely fast rotation reported by Ostro et al. (IAUC 6935) and by Pravec and Šarounová (IAUC 6941).”

SGR 1627–41 M. Feroci, E. Costa, L. Amati, L. Piro, and B. Martino, Istituto di

Astrofisica Spaziale, CNR, Rome; L. Di Ciolo and A. Coletta, BeppoSAX Science Operation Center, Rome; and F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, report: “The Gamma Ray Burst Monitor (GRBM) onboard BeppoSAX has detected bursts from a direction consistent with the new soft γ-ray repeater SGR 1627–41 (IAUC 6944). Two events were triggered on June 17.90091 (consistent with BATSE trigger 6834) and 18.15556 UT. The durations of the events are ∼ 90 and 40 ms, respectively. The spectral indices, assuming a power-law function, are 3.0 ± 0.3 and 2.5 ± 0.3, respectively, in the energy band 40–700 keV. These detections are in agreement with those reported by BATSE and with the harder-energy range of the GRBM. Data analysis is in progress to identify other bursts from the same source.”

NOVA SAGITTARII 1998 Visual magnitude estimates by K. Hornoch, Lelekovice, Czech Rep.:

May 2.07 UT, 10.7; 6.05, 10.6; 10.07, 10.7; 27.04, 11.1; June 1.03, 11.3.

1998 June 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green